At the end of January, Mineral Labs had an inventory of 935 units, which cost S10 per unit to produce. During February, the company produced 1,700 units at a cost of 514 per unit. a. If the firm sold 2,450 units in February, what was the cost of goods sold? (Assume LIFO inventory accounting.) Cost of goods sold b. If the firm sold 2,450 units in February, what was the cost of goods sold? (Assume FIFO inventory accounting.) Cost of goods sold
Answer #1

At the end of January 935 units of inventory @ $10 per unit

In Feb, Company produced 1700 units @ $14 per unit

a.

LIFO Inventory Accounting,

Last Inventory goes out first,

2450 units are sold in Feb,

COGS = 1700(14) + (2450-1700)(10)

COGS = 23800 + 7500

COGS = $31,300

b.

FIFO Inventory Accounting,

Old inventory goes out first,

COGS = 935(10) + (2450-935)(14)

COGS = 9350 + 21210

COGS = $30,560